twelve flat washers, each having an outer diameter of 4" and an inner diameter of 1", are arranged in a 12x16 box in a single layer of four rows having three washers each. if the washers just fit in the box, what is the area, in square inches, of that portion of the base of the box not hidden by washers?
Answer by mananth(16946)

twelve flat washers, each having an outer diameter of 4" and an inner diameter of 1", are arranged in a 12x16 box in a single layer of four rows having three washers each. if the washers just fit in the box, what is the area, in square inches, of that portion of the base of the box not hidden by washers?
...
area of washer = pi*2^2=12.566 in^2
area of hole in the washer = pi*0.5^2=0.785 in^2
Area of box = 12*16=192 in^2
...
192.00-12.566+0.785
=180.22 in^2
...
m.ananth@hotmail.ca